OpenStudy (anonymous):

Matrix help find 1) lA^Tl 2) lA^2l 3) lAA^Tl 4) l2Al 5) lA^(-1)l for: -4 10 5 6

OpenStudy (anonymous):

Determinant of a 2x2 matrix \(\begin{bmatrix}a&b\\c&d\end{bmatrix}\) is \(ad-bc\). \(A^T\) of that same matrix is \(\begin{bmatrix}a&c\\b&d\end{bmatrix}\) . That should be enough info to solve the first few, assuming you know how to multiply matrices. It'd be useful if you could mention specifically where you're getting stuck.

OpenStudy (anonymous):

Here I'm assuming that by using that notation you mean the determinants of each of those matrices. If not, then please clarify.

OpenStudy (maheshmeghwal9):

For a inverse of 2 x 2 matrix if \[[A]=\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\]Then \[[A^{-1}]=\frac{1}{|A|}\left[\begin{matrix}d & -b \\ -c & a \end{matrix}\right]\]

OpenStudy (maheshmeghwal9):

Where |A| = Determinant of [A]

OpenStudy (anonymous):

could you work them out one at the time

OpenStudy (maheshmeghwal9):

Matrix is a vector If 'k' is a scalar multiplied to matrix then; \[[A]=\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]; \space \space k[A]=\left[\begin{matrix}ka & kb \\ kc & kd\end{matrix}\right]\]

OpenStudy (anonymous):

Which part are you having trouble with? You just plug the numbers into the formula.

OpenStudy (maheshmeghwal9):

I think all the theories for ur question is almost completed but where do u have ur acual problem now? I think u must do as @nbouscal says:)

OpenStudy (anonymous):

Here's an example. For your matrix A=\(\begin{bmatrix}-4&10\\5&6\end{bmatrix}\), |A| is \(-4\times6-10\times5=-24-50=-74\)

OpenStudy (maheshmeghwal9):

Here a =-4 b= 10 c= 5 d=6.

OpenStudy (maheshmeghwal9):

now just put the value for the remaining:)

OpenStudy (maheshmeghwal9):

in the formulae; we have given:)

OpenStudy (anonymous):

First calculate each of the matrices, you can check with us that you calculate them right, then after that, calculate their determinants. :)

OpenStudy (zarkon):

you really only need to compute one determinant. The others can be computed using properties of the determinant

OpenStudy (anonymous):

@Zarkon While that is true, I think the exercise was designed to simultaneously give the student practice computing 2x2 determinants and provide a concrete reference point for how the properties of the determinant work.